what is the amount of compound interest on 6,000.00 invested for two years at 9% compounded semiannualy

Answer by checkley77(7064)

P(1+R/N)^TN R=RATE, N=THE NUMBER OF PERIODS PER YEAR, T=NUMBER OF YEARS.
6,000(1+.09/2)^2*2=6,000(1.045)^4=6,000*1.1925=$7,155.11 ANS.